A 0.029 M solution of this ionic compound has an osmotic pressure of 1.50 atm at 25 C has Vans hoff factor that is equal to 3. The vans hoff factor should be express in whole number


1
Expert's answer
2021-02-09T03:46:08-0500

A If FeCl

3

FeCl3 dissociated completely in aqueous solution, it would produce four ions per formula unit [Fe3+(aq) plus 3Cl−(aq)] for an effective concentration of dissolved particles of 4 × 0.0500 M = 0.200 M. The osmotic pressure would be

Π=MRT

=(0.200mol/L)[0.0821(L⋅atm)

/(K⋅mol)](298K)=4.89atm

(13.9.2)

(13.9.2)Π=MRT=(0.200mol/L)[0.0821(L⋅atm)/(K⋅mol)](298K)=4.89atm

B The observed osmotic pressure is only 4.15 atm, presumably due to ion pair formation. The ratio of the observed osmotic pressure to the calculated value is 4.15 atm/4.89 atm = 0.849, which indicates that the solution contains (0.849)(4) = 3.40 particles per mole of FeCl

3

FeCl3 dissolved. Alternatively, we can calculate the observed particle concentration from the osmotic pressure of 4.15 atm:

4.15atm

=M[0.0821(L⋅atm)/(K⋅mol)](298

K)

(13.9.3)

(13.9.3)4.15atm=M[0.0821(L⋅atm)/(K⋅mol)](298K)

0.170mol/L=M

(13.9.4)

(13.9.4)0.170mol/L=M

The ratio of this value to the expected value of 0.200 M is 0.170 M/0.200 M = 0.850, which again gives us (0.850)(4) = 3.40 particles per mole of FeCl

3